# Propeller effect in action: unveiling quenched accretion in the transient X-ray pulsar 4U 0115+63

### v1

@@
The Be/X-ray pulsar 4U 0115+63 underwent a type II outburst in 2023. After the outburst, similar to the outbursts in 2015 and 2017, the source decayed into a quiescent state. Two out of three XMM-Newton observations conducted after the 2023 outburst confirmed the source to be in a low-luminosity state at a level of . X-ray pulsations were detected at 0.277 Hz in both observations with a pulsed fraction exceeding 50%. The power density spectra show no significant low-frequency red noise in both observations, suggesting that the radiation is not driven by accretion. The energy spectra in this state can be described by a single blackbody component, with an emitting area smaller than the typical size of the polar caps during the accretion phase. Based on the timing and spectral properties, we suggest that the propeller effect is active during the quiescent state, resulting in a total quenching of accretion. We discuss possible mechanisms for the generation of pulsations in this regime and consider the scenario of neutron star crust cooling.
@@
Transient X-ray pulsars (XRPs) in Be X-ray binaries (BeXRPs) host a highly magnetized neutron star (NS) with a magnetic field of the order – G. BeXRPs emits X-rays when accreting matter from a decretion disk that episodically forms around a rapidly rotating Be-type companion star. As the accreted material approaches the NS magnetosphere, it couples to the magnetic field lines and is funneled onto the magnetic poles. This process releases gravitational potential energy, which is radiated as X-rays. The anisotropic nature of the accretion flow leads to pulsed X-ray emission modulated by the NS rotation (for a review, see Mushtukov & Tsygankov, 2024). These systems usually have significantly eccentric orbits. Near periastron, the increase in mass accretion rate can trigger a so-called type I outburst, with X-ray luminosities reaching up to . In other cases, BeXRPs can reach luminosities above the Eddington limit for a NS during giant (type II) outbursts. The physical mechanism behind type II outbursts remains uncertain (see review by Reig, 2011).
@@
The temporal evolution of the observed X-ray flux from transient XRPs is governed to a large extent by the interaction of matter with the magnetic field. For instance, for the systems with fast-spinning NSs (typical spin period ), at the end of the outburst, when the mass accretion rate decreases to a threshold value, the accretion is expected to be inhibited by the centrifugal barrier – the so-called “propeller regime” (Illarionov & Sunyaev, 1975). The threshold mass accretion rate corresponding to the transition to the propeller regime corresponds to the equality of the magnetospheric and co-rotation radii of the NS, and can be expressed as (e.g., Stella et al., 1986):
@@
where is the magnetic field strength in units of , is the spin period in seconds, is the NS mass in units of , is the radius in , and is a geometric factor of order unity. Below this luminosity threshold, accretion onto the NS is expected to cease abruptly, resulting in a sharp decline in observed flux and a transition to a low-luminosity quiescent state. In several XRPs, dramatic luminosity evolutions have been observed in the final stages of their outbursts and are considered to result from the propeller effect, e.g., SXP 4.78 (Semena et al., 2019), GRO J175027 (Lutovinov et al., 2019), 4U 0115+63, and V 0332+53 (Stella et al., 1986; Campana et al., 2001; Tsygankov et al., 2016).
@@
At the same time, most transient XRPs observed in quiescence exhibit a non-zero flux in the X-ray band, with a typical luminosity of (Tsygankov et al., 2017b). Moreover, some sources still display pulsations during this state (e.g., 4U 1145619. Mukherjee & Paul 2005; 1A 1118615, Rutledge et al. 2007; 1A 0535+262, Mukherjee & Paul 2005; 4U 0115+63, Rouco Escorial et al. 2017, 2020; see also Table 1 in Tsygankov et al. 2017b for a complete source list). The reason for this phenomenon remains unknown, with two main hypotheses proposed. The first is the leakage of matter through the centrifugal barrier along magnetic field lines, leading to the formation of hot spots (see, e.g., Elsner & Lamb, 1977; Ikhsanov, 2001; Lii et al., 2014; Orlandini et al., 2004; Mukherjee & Paul, 2005). Another possible mechanism is thermal emission from a cooling, accretion-heated NS, in which more heat from the solid crustal layer is conducted to the magnetic poles due to anisotropic heat transport governed by the magnetic field lines (Rouco Escorial et al., 2017).
@@
4U 0115+63 is a typical transient BeXRP that was discovered by the Uhuru satellite (Giacconi et al., 1972) and identified as having a B0.2Ve spectral type companion star (Johns et al., 1978). The NS has a magnetic field strength of measured from the cyclotron line energy (Heindl et al., 2000) and a spin period of (Cominsky et al., 1978). It orbits its companion every 24.3 d with an eccentricity of approximately 0.34 (Rappaport et al., 1978). The distance to the source is estimated to be , based on the Gaia data (Neumann et al., 2023).
@@
In this paper, we perform timing and spectral analysis of the deep XMM-Newton observations of 4U 0115+63 obtained after the 2023 giant outburst to constrain the origin of its pulsating emission in the quiescent state.
@@ 2 Observations and data reduction
In March 2023, 4U 0115+63 exhibited a new type II outburst, followed by a decay into a low-luminosity (quiescent) state. XMM-Newton observed the source multiple times, and here we use the suitable data from the XMM European Photon Imaging Camera (EPIC), specifically the EPIC-PN and EPIC-MOS detectors. Observations were conducted on 2023 July 13 (MJD 60138, Obs1), August 4 (MJD 60160, Obs2), and September 11 (MJD 60198, Obs3). We also include an observation conducted on 2016 February 17 (MJD 57435; Obs2016), when the source was in the quiescence (see Table 1 and Fig. 1 for details).
@@
Following the official SAS Science threads,222https://www.cosmos.esa.int/web/xmm-newton/sas-threads the data reduction was performed by using the Science Analysis System (SAS v22.1.0)333https://www.cosmos.esa.int/web/xmm-newton/sas software with the latest calibration datasets. We used the epproc and emproc tasks to produce calibrated and concatenated event lists from the EPIC-PN (hereafter PN) and EPIC-MOS (hereafter MOS) cameras, respectively. To remove high background flaring activity from the observations, the single-event (PATTERN==0) light curves in the 10–12 keV band for the PN camera and 10 keV for the MOS cameras were extracted; then we used the threshold rates for a steady low background (see Table 1) to get the good time intervals (GTIs) of each observation. Due to the very low count rates of the source during the XMM-Newton observations, Obs1, Obs2, and Obs2016 were not affected by pile-up. For these observations, the source region was identified as a circular region with a 20″ radius, and the background was extracted from a nearby, source-free circular region with a 50″ radius on the same CCD chip. For Obs3, to mitigate pile-up effects, we used an annular source region centered on the source position, with an inner radius of 10″ and an outer radius of 20″. We extracted the single and double events (PATTERN 4) for the PN camera and the events between single and quadruple (PATTERN 12) for the MOS cameras. The barycentric correction was performed by using barycen task.
@@
For the timing analysis, we used the epiclccorr task to obtain net light curves. We used the powspec tool to get the power density spectra (PDSs) and folded the pulse profiles using the efold tool from the HEASoft v6.35.1 package.444https://heasarc.gsfc.nasa.gov/docs/software/lheasoft The background spectra were scaled using the backscale task. The response matrix and the ancillary response files were generated using rmfgen and arfgen, respectively. We used the epicspeccombine task to combine the spectra of three EPIC cameras in order to improve statistics. Due to the very few counts in the spectra, we rebinned the spectra with grppha tool to make sure each bin has a minimum count of 5. The spectral analysis was performed by using the X-ray spectral fitting package (xspec) version 12.15.0 (Arnaud, 1996). We used C-statistics for the statistics of fitting. All uncertainties in this paper for spectral analysis correspond to a confidence level of 90%, estimated by running Monte Carlo Markov Chains with a length of 20,000 with a burn-in of 10,000 using the Goodman-Weare algorithm.
@@ 3.1 Timing analysis
In accreting systems like XRPs, accretion is an intrinsically noisy process, where fluctuations in the mass accretion rate are driven by random variations in the viscosity within the accretion disk (Lyubarskii, 1997). Therefore, if accretion is the primary mechanism responsible for the observed X-ray emission, a feasible way to validate this process is by studying the aperiodic variability in the light curves (e.g., Doroshenko et al., 2020). Specifically, the authors demonstrated a comparison of PDS between accreting pulsars and non-accreting systems (like magnetars and radio pulsars), showing that low-frequency noise dominates the PDS of accreting systems. They attributed these differences to the presence or absence of ongoing accretion onto the NS.
To test for the presence of accretion in 4U 0115+63 during the different states observed with XMM-Newton, we generated PDS using the powspec tool from the HEASoft package, as shown in Fig. 3. We note that the XMM-Newton observation from 2016 (Obs2016) shows an averaged count rate similar to that of Obs1 and Obs2 (Rouco Escorial et al., 2020). Therefore, we included this observation in the timing and spectral analysis to improve the statistics. As clearly seen in Fig. 3, the red noise dominates the low-frequency part of the PDS in Obs3. This observation was performed when the NS was near periastron passage (see Fig. 1), a phase where the system is expected to experience an increased accretion rate, potentially leading to a type I outburst. In contrast, the PDS of the merged light curves from Obs1, Obs2, and Obs2016 remains flat at a level of at low frequencies, which is characteristic of the white noise. We fitted this PDS with a constant value of 2, yielding a reduced chi-square of with 511 degrees of freedom. At the same time, peaks in the PDS corresponding to the spin frequency are significantly detected at 0.277 Hz in both states. We fitted both peaks with a gaussian function, resulting in a standard deviation of Hz for the quiescent observations, whereas the peak in Obs3 appears broadened with a higher standard deviation of Hz, likely due to the interplay between periodic and aperiodic variability.
The pulse profiles of 4U 0115+63 obtained when the source was in the quiescent (Obs1, Obs2, and Obs2016) and accreting states (Obs3) are shown in Fig. 4. For the quiescence, the pulse profiles are very similar, consisting of a sinusoidal-like peak accompanied by a weak harmonic, consistent with the findings of Rouco Escorial et al. (2020). The averaged pulse profile was fitted with a combination of a sinusoid and its first harmonic, revealing an amplitude ratio between the fundamental and harmonic of approximately 3. We quantified the pulse amplitude using the fractional root-mean-square pulsed fraction (), which is defined as:
@@ 3.2 Spectral analysis
According to observational studies (Tsygankov et al., 2016; Wijnands & Degenaar, 2016; Rouco Escorial et al., 2017, 2020), the continuum emission of 4U 0115+63 in the quiescent states can be described by a blackbody model. We adopted a spectral model consisting of a blackbody component (bbodyrad in xspec) modified by interstellar photoelectric absorption (tbabs in xspec). The hydrogen column density, , was fixed at following Wijnands & Degenaar (2016). Element abundances were set to wilm (Wilms et al., 2000) and photoelectric cross-sections to vern (Verner et al., 1996). As shown in Table 6, the fitting results shows an acceptable goodness-of-fit statistics: C-statistic/dof=86.1/86, 28.1/37, and 229.6/202 for Obs1, Obs2, and Obs2016, respectively. The best-fit blackbody temperatures () for these three spectra are , , and keV respectively. Assuming isotropic emission and a distance of 5.0 kpc, the corresponding emission region radii, derived from the model normalization, are , , and km, respectively. These radii are significantly smaller than the canonical NS radius, suggesting that the X-ray emission originates from localized hot spots at the magnetic poles of the NS. For Obs3, the spectrum can not be described by a single bbodyrad. We therefore fitted the spectrum using an absorbed power-law model, i.e., tbabs*(powerlaw), which provides a good fit with /dof= 734.81/732. The resulting parameters indicate a hard photon index of , which is typical of the accreting state.
@@
Similar to the timing analysis (e.g., the presence of red noise), the spectral shape of an XRP’s emission may provide strong evidence either for ongoing accretion or for its absence. Notably, Tsygankov et al. (2019b) demonstrated that even at low mass accretion rates, the X-ray spectrum of XRPs cannot be adequately described by a simple blackbody model. In practice, at low luminosities (), XRP spectra typically exhibit a characteristic “two-hump” shape. This feature has been reported in several sources, including X Persei (Doroshenko et al., 2012), GX 3041 (Tsygankov et al., 2019b), 1A 0535+262 (Tsygankov et al., 2019a), GRO J100857 (Lutovinov et al., 2021), and 2RXP J130159.6635806 (Salganik et al., 2025). The hard spectral component is thought to arise partly from cyclotron radiation, originating from interaction of the accretion flow with the NS atmosphere, Comptonized in its upper hot layers (Nelson et al., 1995; Di Salvo et al., 1998; Coburn et al., 2001; Tsygankov et al., 2019a; Mushtukov et al., 2021; Sokolova-Lapa et al., 2021).
Mushtukov et al. (2021) and Sokolova-Lapa et al. (2021) modeled this scenario and successfully reproduced the overall shape of the low-luminosity spectra of A 0535+262 and GX 3041. Both models predict that the position of the hard component coincides with the cyclotron energy of the pulsar. To test for the presence of a hard emission component related to the accretion process during the quiescent state, we added a gaussian component to our spectral model (i.e. const*tbabs*bbodyrad), fixing its central energy at 12 keV, corresponding to the fundamental cyclotron line of 4U 0115+63 (e.g, White et al. 1983). Following the suggestion by Tsygankov et al. (2019b) that the hard component typically contributes about half the flux of the soft component, we constrained the flux of the gaussian component in our fitting procedure to be half that of the blackbody component. However, the best-fit results gave a width of the gaussian close to zero. This suggests that the hard component is not necessary for the spectral fitting. We further fixed its width at 10 keV to enable a comparison between fits with and without the hard component, resulting a goodness-of-fit of C-statistic/dof=392.2/325. The negative value of indicates that the two-component model provides a worse fit to the spectrum compared to the single-blackbody model. Therefore, the presence of a hard component in the quiescence spectrum can be tentatively ruled out. The corresponding fits are shown in Fig. 5. However, we note that the limited energy coverage of the available data restricts one to fully characterize the spectra above 10 keV, the spectral results should be interpreted with caution. Nevertheless, due to the low cyclotron line energy of 4U 0115+63, future broadband observations of this source would be valuable for understanding the radiation mechanisms in the low-luminosity state.
@@
- | - | - | 1.14 (fixed) | 1.14 (fixed) | 1.14 (fixed) | 1.14 (fixed) | ||
- | - | - | 1 (fixed) | 1 (fixed) | 1 (fixed) | 1 (fixed) | 1 (fixed) | |
- | - | - | 0.73 (fixed) | 0.73 (fixed) | 0.73 (fixed) | 0.73 (fixed) | ||

### v2

@@
The Be/X-ray pulsar 4U 0115+63 underwent a type II outburst in 2023. After the outburst, similar to the outbursts in 2015 and 2017, the source decayed into a quiescent state. Two out of three XMM-Newton observations conducted after the 2023 outburst confirmed the source to be in a low-luminosity state at a level of . X-ray pulsations were detected at 0.277 Hz in both observations with a pulsed fraction exceeding 50%. The power density spectra show no significant low-frequency red noise in either observation, suggesting that the radiation is not driven by accretion. The energy spectra in this state can be described by a single blackbody component, with an emitting area smaller than the typical size of the polar caps during the accretion phase. Based on the timing and spectral properties, we suggest that the propeller effect is active during the quiescent state, resulting in a total quenching of accretion. We discuss possible mechanisms for the generation of pulsations in this regime and consider the scenario of neutron star crust cooling.
@@
Transient X-ray pulsars (XRPs) in Be X-ray binaries (BeXRPs) host a highly magnetized neutron star (NS) with a magnetic field on the order of – G. BeXRPs emit X-rays when accreting matter from a decretion disk that episodically forms around a rapidly rotating Be-type companion star. As the accreted material approaches the NS magnetosphere, it couples to the magnetic field lines and is funneled onto the magnetic poles. This process releases gravitational potential energy, which is radiated as X-rays. The anisotropic nature of the accretion flow leads to pulsed X-ray emission modulated by the NS rotation (for a review, see Mushtukov & Tsygankov, 2024). These systems usually have significantly eccentric orbits. Near periastron, the increase in the mass accretion rate can trigger a so-called type I outburst, with X-ray luminosities reaching up to . In other cases, BeXRPs can reach luminosities above the Eddington limit for a NS during giant (type II) outbursts. The physical mechanism behind type II outbursts remains uncertain (see the review by Reig, 2011).
@@
The temporal evolution of the observed X-ray flux from transient XRPs is governed to a large extent by the interaction of matter with the magnetic field. For instance, for the systems with fast-spinning NSs (typical spin period ), at the end of the outburst, when the mass accretion rate decreases to a threshold value, the accretion is expected to be inhibited by the centrifugal barrier – the so-called propeller regime (Illarionov & Sunyaev, 1975). The threshold mass accretion rate corresponding to the transition to the propeller regime corresponds to the equality of the magnetospheric and corotation radii of the NS and can be expressed as (e.g., Stella et al., 1986)
@@
where is the magnetic field strength in units of , is the spin period in seconds, is the NS mass in units of , is the radius in , and is a geometric factor of order unity. Below this luminosity threshold, accretion onto the NS is expected to cease abruptly, resulting in a sharp decline in observed flux and a transition to a low-luminosity quiescent state. In several XRPs, dramatic luminosity evolutions have been observed in the final stages of their outbursts and are considered to result from the propeller effect, for example SXP 4.78 (Semena et al., 2019), GRO J175027 (Lutovinov et al., 2019), 4U 0115+63, and V 0332+53 (Stella et al., 1986; Campana et al., 2001; Tsygankov et al., 2016).
@@
At the same time, most transient XRPs observed in quiescence exhibit a nonzero flux in the X-ray band, with a typical luminosity of (Tsygankov et al., 2017b). Moreover, some sources still display pulsations during this state (e.g., 4U 1145619, Mukherjee & Paul 2005; 1A 1118615, Rutledge et al. 2007; 1A 0535+262, Mukherjee & Paul 2005; and 4U 0115+63, Rouco Escorial et al. 2017, 2020; see also Table 1 in Tsygankov et al. 2017b for a complete source list). The reason for this phenomenon remains unknown, with two main hypotheses proposed. The first is the leakage of matter through the centrifugal barrier along magnetic field lines, which leads to the formation of hot spots (see, e.g., Elsner & Lamb, 1977; Ikhsanov, 2001; Lii et al., 2014; Orlandini et al., 2004; Mukherjee & Paul, 2005). The second is thermal emission from a cooling, accretion-heated NS, in which more heat from the solid crustal layer is conducted to the magnetic poles due to anisotropic heat transport governed by the magnetic field lines (Rouco Escorial et al., 2017).
@@
4U 0115+63is a typical transient BeXRP that was discovered by the Uhuru satellite (Giacconi et al., 1972) and identified as having a B0.2Ve spectral type companion star (Johns et al., 1978). The NS has a magnetic field strength as measured from the cyclotron line energy (Heindl et al., 2000) and a spin period (Cominsky et al., 1978). It orbits its companion every 24.3 d with an eccentricity of approximately 0.34 (Rappaport et al., 1978). The distance to the source is estimated to be , based on the Gaia data (Neumann et al., 2023).
@@
In this study we performed timing and spectral analysis of the deep XMM-Newton observations of 4U 0115+63 obtained after the 2023 giant outburst to constrain the origin of its pulsating emission in the quiescent state.
@@ 2 Observations and data reduction
In March 2023, 4U 0115+63 exhibited a new type II outburst, followed by a decay into a low-luminosity (quiescent) state. XMM-Newton observed the source multiple times, and we used the suitable data from the XMM European Photon Imaging Camera (EPIC), specifically the EPIC-PN (hereafter PN) and EPIC-MOS (hereafter MOS) detectors. Observations were conducted on 2023 July 13 (MJD 60138, hereafter Obs1), August 4 (MJD 60160, Obs2), and September 11 (MJD 60198, Obs3). We also included an observation conducted on 2016 February 17 (MJD 57435; Obs2016), when the source was in the quiescence (see Table 1 and Fig. 1 for details).
@@
Following the official SAS Science threads,222https://www.cosmos.esa.int/web/xmm-newton/sas-threads the data reduction was performed by using the Science Analysis System (SAS v22.1.0)333https://www.cosmos.esa.int/web/xmm-newton/sas software with the latest calibration datasets. We used the epproc and emproc tasks to produce calibrated and concatenated event lists from the PN and MOS cameras, respectively. To remove high background flaring activity from the observations, the single-event (PATTERN==0) light curves in the 10–12 keV band for the PN camera and 10 keV for the MOS cameras were extracted; then we used the threshold rates for a steady low background (see Table 1) to get the good time intervals of each observation. Due to the very low count rates of the source during the XMM-Newton observations, Obs1, Obs2, and Obs2016 were not affected by pile-up. For these observations, the source region was identified as a circular region with a 20″ radius, and the background was extracted from a nearby, source-free circular region with a 50″ radius on the same CCD chip. For Obs3, to mitigate pile-up effects, we used an annular source region centered on the source position, with an inner radius of 10″ and an outer radius of 20″. We extracted the single and double events (PATTERN 4) for the PN camera and the events between single and quadruple (PATTERN 12) for the MOS cameras. The barycentric correction was performed by using barycen task.
@@
For the timing analysis, we used the epiclccorr task to obtain net light curves. We used the powspec tool to get the power density spectra (PDSs) and folded the pulse profiles using the efold tool from the HEASoft v6.35.1 package.444https://heasarc.gsfc.nasa.gov/docs/software/lheasoft The background spectra were scaled using the backscale task. The response matrix and the ancillary response files were generated using rmfgen and arfgen, respectively. We used the epicspeccombine task to combine the spectra of three EPIC cameras in order to improve statistics. Due to the very few counts in the spectra, we re-binned the spectra with grppha tool to make sure each bin has a minimum count of 5. The spectral analysis was performed by using the X-ray spectral fitting package (xspec) version 12.15.0 (Arnaud, 1996). We used C-statistics for the statistics of fitting. All uncertainties in this paper for spectral analysis correspond to a confidence level of 90%, estimated by running Monte Carlo Markov chains with a length of 20,000 with a burn-in of 10,000 using the Goodman-Weare algorithm.
@@ 3.1 Timing analysis
In accreting systems like XRPs, accretion is an intrinsically noisy process, where fluctuations in the mass accretion rate are driven by random variations in the viscosity within the accretion disk (Lyubarskii, 1997). Therefore, if accretion is the primary mechanism responsible for the observed X-ray emission, a feasible way to validate this process is by studying the aperiodic variability in the light curves (e.g., Doroshenko et al., 2020). Specifically, the authors compared the PDSs of accreting pulsars and non-accreting systems (such as magnetars and radio pulsars), showing that low-frequency noise dominates the PDSs of accreting systems. They attributed these differences to the presence or absence of ongoing accretion onto the NS.
To test for the presence of accretion in 4U 0115+63 during the different states observed with XMM-Newton, we generated PDSs using the powspec tool from the HEASoft package, as shown in Fig. 3. We note that the XMM-Newton observation from 2016 (Obs2016) shows an averaged count rate similar to that of Obs1 and Obs2 (Rouco Escorial et al., 2020). Therefore, we included this observation in the timing and spectral analysis to improve the statistics. As clearly seen in Fig. 3, the red noise dominates the low-frequency part of the PDS in Obs3. This observation was performed when the NS was near periastron passage (see Fig. 1), a phase where the system is expected to experience an increased accretion rate, potentially leading to a type I outburst. In contrast, the PDS of the merged light curves from Obs1, Obs2, and Obs2016 remains flat at a level of at low frequencies, which is characteristic of the white noise. We fitted this PDS with a constant value of 2, yielding a reduced chi-square of with 511 degrees of freedom. At the same time, peaks in the PDS corresponding to the spin frequency are significantly detected at 0.277 Hz in both states. We fitted both peaks with a Gaussian function, resulting in a standard deviation of Hz for the quiescent observations, whereas the peak in Obs3 appears broadened with a higher standard deviation of Hz, likely due to the interplay between periodic and aperiodic variability.
The pulse profiles of 4U 0115+63 obtained when the source was in the quiescent (Obs1, Obs2, and Obs2016) and accreting states (Obs3) are shown in Fig. 4. For the quiescence, the pulse profiles are very similar, consisting of a sinusoidal-like peak accompanied by a weak harmonic, consistent with the findings of Rouco Escorial et al. (2020). The averaged pulse profile was fitted with a combination of a sinusoid and its first harmonic, revealing an amplitude ratio between the fundamental and harmonic of approximately 3. We quantified the pulse amplitude using the fractional root-mean-square pulsed fraction (), which is defined as
@@ 3.2 Spectral analysis
According to observational studies (Tsygankov et al., 2016; Wijnands & Degenaar, 2016; Rouco Escorial et al., 2017, 2020), the continuum emission of 4U 0115+63 in the quiescent states can be described by a blackbody model. We adopted a spectral model consisting of a blackbody component (bbodyrad in xspec) modified by interstellar photoelectric absorption (tbabs in xspec). The hydrogen column density, , was fixed at following Wijnands & Degenaar (2016). Element abundances were set to wilm (Wilms et al., 2000) and photoelectric cross-sections to vern (Verner et al., 1996). As shown in Table 6, the fitting results shows an acceptable goodness-of-fit statistics: C-statistic/dof=86.1/86, 28.1/37, and 229.6/202 for Obs1, Obs2, and Obs2016, respectively. The best-fit blackbody temperatures () for these three spectra are , , and keV, respectively. Assuming isotropic emission and a distance of 5.0 kpc, the corresponding emission region radii, derived from the model normalization, are , , and km, respectively. These radii are significantly smaller than the canonical NS radius, suggesting that the X-ray emission originates from localized hot spots at the magnetic poles of the NS. For Obs3, the spectrum cannot be described by a single bbodyrad. We therefore fitted the spectrum using an absorbed power-law model, i.e., tbabs*(powerlaw), which provides a good fit with /dof= 734.81/732. The resulting parameters indicate a hard photon index of , which is typical of the accreting state.
@@
Similar to the timing analysis (e.g., the presence of red noise), the spectral shape of an XRP’s emission may provide strong evidence of either ongoing accretion or its absence. Notably, Tsygankov et al. (2019b) demonstrated that even at low mass accretion rates, the X-ray spectrum of XRPs cannot be adequately described by a simple blackbody model. In practice, at low luminosities (), XRP spectra typically exhibit a characteristic “two-hump” shape. This feature has been reported in several sources, including X Persei (Doroshenko et al., 2012), GX 3041 (Tsygankov et al., 2019b), 1A 0535+262 (Tsygankov et al., 2019a), GRO J100857 (Lutovinov et al., 2021), and 2RXP J130159.6635806 (Salganik et al., 2025). The hard spectral component is thought to arise partly from cyclotron radiation, originating from interaction of the accretion flow with the NS atmosphere, Comptonized in its upper hot layers (Nelson et al., 1995; Di Salvo et al., 1998; Coburn et al., 2001; Tsygankov et al., 2019a; Mushtukov et al., 2021; Sokolova-Lapa et al., 2021).
Mushtukov et al. (2021) and Sokolova-Lapa et al. (2021) modeled this scenario and successfully reproduced the overall shape of the low-luminosity spectra of A 0535+262 and GX 3041. Both models predict that the position of the hard component coincides with the cyclotron energy of the pulsar. To test for the presence of a hard emission component related to the accretion process during the quiescent state, we added a gaussian component to our spectral model (i.e., const*tbabs*bbodyrad), fixing its central energy at 12 keV, corresponding to the fundamental cyclotron line of 4U 0115+63 (e.g., White et al. 1983). Following the suggestion by Tsygankov et al. (2019b) that the hard component typically contributes about half the flux of the soft component, we constrained the flux of the gaussian component in our fitting procedure to be half that of the blackbody component. However, the best-fit results gave a width of the gaussian close to zero. This suggests that the hard component is not necessary for the spectral fitting. We further fixed its width at 10 keV to enable a comparison between fits with and without the hard component, resulting a goodness-of-fit of C-statistic/dof=392.2/325. The negative value of indicates that the two-component model provides a worse fit to the spectrum compared to the single-blackbody model. Therefore, the presence of a hard component in the quiescence spectrum can be tentatively ruled out. The corresponding fits are shown in Fig. 5. However, we note that the limited energy coverage of the available data restricts one to fully characterize the spectra above 10 keV, the spectral results should be interpreted with caution. Nevertheless, due to the low cyclotron line energy of 4U 0115+63, future broadband observations of this source would be valuable for understanding the radiation mechanisms in the low-luminosity state.
@@
| - | - | - | 1.14 (fixed) | 1.14 (fixed) | 1.14 (fixed) | 1.14 (fixed) | ||
| - | - | - | 1 (fixed) | 1 (fixed) | 1 (fixed) | 1 (fixed) | 1 (fixed) | |
| - | - | - | 0.73 (fixed) | 0.73 (fixed) | 0.73 (fixed) | 0.73 (fixed) | ||